you have one hundred coins in your coin jar, consisting of nickels, dimes and quarters. the total value of the coins is &11.80 and you have 10 more nickels than dimes. How many of each coin do you have?
x=dimes
x +10= nickels
90-2x= quarters

Answer by Alan3354(30993)

you have one hundred coins in your coin jar, consisting of nickels, dimes and quarters. the total value of the coins is &11.80 and you have 10 more nickels than dimes. How many of each coin do you have?
x=dimes
x +10= nickels
90-2x= quarters
-----------------
What you have is correct.
Then add:
10*x + 5*(x+10) + 25*(90-2x) = 1180
You can do it from here, I'm sure, but...
10x + 5x+50 + 2250-50x = 1180
-35x = -1120
x = 32 dimes (= $3.20)
42 = nickels (= $2.10)
26 = quarters (= $6.50)
100 coins------$11.80